OpenStudy (anonymous):

Find the domain of the function. f(x)= sqrt*6-x

OpenStudy (anonymous):

@kl0723

OpenStudy (anonymous):

B. (-∞,sqrt*6 ) (sqrt*6 , ∞) C. (-∞,sqrt*6 ] D. (-∞, 6]

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

c

OpenStudy (anonymous):

you cant find the square root of a negative and anything bigger than six would do it when plugged into the x value

OpenStudy (triciaal):

you know the domain is the values of x that will satisfy f(x)

OpenStudy (anonymous):

Complete the square and write the equation in standard form. Then give the center and radius of the circle x^2 + y^2 - 10x - 8y + 29 = 0

OpenStudy (anonymous):

do ur homework yo

OpenStudy (anonymous):

I have but I got these two questions wrong I just need help on this last question notice theres not 4 answers because, I already got it wrong.

OpenStudy (anonymous):

B. (x - 5)2 +(y - 4)2 = 12 (-5, -4), r = 2 sqrt*3 c.(x - 5)2 +(y - 4)2 = 12 (5, 4), r = 2 sqrt*3 D. (x + 5)2 +(y + 4)2 = 12 (-5, -4), r = 2 sqrt*3

OpenStudy (triciaal):

does this mean you had choice a and it was wrong?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

c?

OpenStudy (triciaal):

I don't see how the other person above had the square root of a negative I would take the + and - of rt 6 got it !!